40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8165     Entry time: Tue Feb 26 01:49:27 2013     In reply to: 8162     Reply to this: 8166
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI locked 

[Jenne, Yuta]

We began the evening, after alignment of all optics was good (arms flashing, PRC flashing, assumed SRM last saved alignment was okay), centering all oplevs and aligning beam onto AS55, REFL11, REFL55 and REFL33 and POPDC.

After a quick check to make sure that the input pointing was still okay for Yarm (TRY was 0.88 when we began PRMI work, which we called okay), we aligned and locked the Michelson with AS55Q.  We were able to use a gain as large (abs val) as -15 before the loop started oscillating.  (ETMs, PRM, SRM all misaligned during this).  We measured the UGF of the MICH loop to be 170Hz, with phase margin of 40 degrees.

We then restored the PRM, and tweaked the pointing until the PRC beam at AS overlapped the MICH beam. 

We then started playing with locking.  We were not very successful in using REFL 11 or REFL 55 (I for both, although we also tried 11Q just for kicks).  We then switched to using REFL33I, and had success!!  We are reliably able to lock to the "sideband", and not so reliably lock to the carrier (by flipping the sign of the PRCL loop gain). I say "sideband" with quotes, since we aren't sure that it is the sideband.  We are, however, confident that it is locking, and it's certainly not locked to the carrier.  Videos are at the bottom of the entry.

A list of some values:

Camera state MICH gain PRCL gain POP DC AS DC REFL DC
AS bright, POP bright -0.045 -4.000 1000 +/- 200 4 +/- 0.5  480
AS dark, POP bright  -0.045  +4.000  300 +/- 20  0.03 +/- 0.02  552 +/- 2
AS, POP flashing  loop off  loop off  flash 25,000 or more  ~5  ~550

 

Other notes:  We changed AS55's demod phase back to 24.5, from it's atmosphere half-cavity value.  The change from the original value was recorded in elog 8030.

We changed REFL11's demod phase back to 150, which is the value that it was when we had PRMI locked on ~July 10th, 2012.  (We looked up the burt snapshot to check).

 

FI back, upper right is POP, lower left is REFL, lower right is AS.  It seems as though we may need to redo coil balancing now that we're at vacuum / with the current OSEM values.

ELOG V3.1.3-